a police jeep is chasing a culprit going on a motorbike . the motorbike crosses a turning ata 72 km / h the jeep follows it at 90 km/ h , crossing the turnings 10 s later than the bike . assuming that they travel at constant speed s , how far from the turnings will yhe jeep catch up with the bike ?

Grade:11

1 Answers

Saransh Anand
7 years ago
VP = 90 km/h = 25 m/s.
VC = 72 km/h = 20 m/s.
 
Time taken by the culprit to reach B from A= 10s
Distance converted by culprit  = v x t = 20 x 10 = 200 m.
 
At time t = 10 sec the police jeep is 200 m behind the
culprit.

Time = s/v = 200 / 5 = 40 s. (Relative velocity is considered).
In 40 s the police jeep will move from A to a distance S, where
S = v x t = 25 x 40 = 1000 m = 1.0 km away.

Thus the jeep will catch up with the bike, 1 km far from the turning.